OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 120r^3 square inches. Each tile has an area of sqrt(16r^9). Use the given information to write and simplify an expression for the total number of tiles used.

OpenStudy (anonymous):

120r^3/sqrt(16r^9) 120r^3/4r^3 30 Trevor must have 30 tiles in order to tile his bathroom floor. I was told that my second line was incorrect

OpenStudy (anonymous):

yes it is wrong \[\sqrt{16t^9}=\sqrt{4^2\times r^8\times r}=\sqrt{4^2}\sqrt{r^8}\sqrt{r}=4r^4\sqrt{r}\]

OpenStudy (anonymous):

So how would I divide it?

OpenStudy (anonymous):

\[\frac{120r^3}{4r^4\sqrt{r}}=\frac{30}{r\sqrt{r}}\]

OpenStudy (anonymous):

if you are not allowed to have radicals in the denominator, i guess you could write \[\frac{30\sqrt{r}}{r^2}\]